OpenStudy (christos):

Find the period of the function f(x)=2cos(4x+3π)

OpenStudy (anonymous):

In this case, because it's a circular function, the domain will go on forever unless you restrict it. So, in this case, going by the information, the domain is infinity. The range will be [-3, 1], because of the 2 in front of the cos. The period is how long the function takes to end up back where it started, or to make a full cycle. The period would be pi/2 for this.

OpenStudy (anonymous):

\[T=\frac{2\pi}{b}=\frac{2\pi}{4}=\frac{\pi}{2}\]

OpenStudy (anonymous):

for \[f(x)=asin(bx-c)+d\]

OpenStudy (anonymous):

T=2πb=2π4=π2 this is just simple form of what i said

OpenStudy (anonymous):

yeah just show the work... some people get mad about random answers as it doesn't really show the person that needs help

OpenStudy (christos):

@Outkast3r09 where did you make use of the number 2 from the function at "2cos" and if yes where?

OpenStudy (anonymous):

nope just used b as a is just the amplitude, it tells you how highs and lows (max and mins) of the sine function whereas the period is how many rads it takes to do a complete cycle

OpenStudy (christos):

where is the "B" in this specific function?

OpenStudy (anonymous):

\[f(x)=acos(bx-c)+d\]

OpenStudy (anonymous):

T=2πb=2π4=π2

OpenStudy (anonymous):

it's the coefficient infront of x

OpenStudy (christos):

ok and something last. What did you mean in your previews respond when you said " for f(x)=asin(bx−c)+d" ?

OpenStudy (christos):

I mean.... Every time I need to find a period I use this formula?

OpenStudy (christos):

T = blabla?

OpenStudy (anonymous):

t = F(x)

OpenStudy (christos):

@Outkast3r09 Do I use this every time I wanna find a period? T=2πb=2π4=π2

OpenStudy (anonymous):

yes mam

OpenStudy (anonymous):

yes however the top will change. If you read Zoodude's explanation or mine . Period is the time it takes for the function to complete a full cycle. Sine and Cos both reach a complete cycle at 2 pi, whereas tan is at pi. the coefficient b affects the rads for completion of function

OpenStudy (anonymous):

that is why T is the relation between normal period and the shrinking / stretching of the graph

OpenStudy (christos):

so for tan and cot is T=π/b

OpenStudy (christos):

?

OpenStudy (anonymous):

yes

OpenStudy (christos):

what about csc and sec?